%age of Se in peroxidase anhydrous enzyme is 0.5% by weight then molecular weight of peroxidase anhydrous enzyme.

Grade:12th pass

1 Answers

Arun
4 years ago
let molecular weight of enzyme be x
weight of Se in enzyme = 0.5% of x 
 = 0.5 x / 100
at. wt. of Se = 78.4 (which is given in the question)
so,
78.4=0.5 x / 100
78.4 * 100 = 0.5 x 
7840 = 0.5 x 
x = 7840 / 0.5
x = 7840 * 10 / 5
therefore , x = 15680 or 1.5680 * 10^4